What is the formula for solving this: "If you draw 4 M&Ms from a bag containing 4 red, 4 yellow, 4 green, and 4 brown ones, what is the probability that you will get 1 of each color?" It boggles my mind. Thanks in advance!

Found 2 solutions by Greenfinch, scott8148:
Answer by Greenfinch(383)

1st pick p = 1 (your are going to get a sweet and it doesn't matter which colour)
2nd pick p= 12/15 anything but 1st colour (of which 3 are left)
3rd pick p = 8/14 two colours left out of 4
4th p = 4/13
So probability is = 1 x 12/15 x 8/14 x 4/13 = 64/(5 x 7 x 13)= 64/455

Answer by scott8148(6628)

number of ways to draw one of each ___ 4 * 4 * 4 * 4 = 256

number of ways to draw 4 from 16 ___ 16C4 = (16 * 15 * 14 * 13) / (4 * 3 * 2 * 1) = 1820

probability is ___ 256 / 1820 ___ or .141 (approx)